OpenStudy (anonymous):

find the vertex,focus and directrix and the correct graph of the equation 8(x+3)=(y+2)^2

OpenStudy (campbell_st):

rewrite the equation 8(x + 3) = (y+2)^2 general form for a parabola about the y-axis is 4a( x -h) = (y-k)^2 (h, k) is the vertex, x = -a is the equation of the directrix and (h+a, k) is the focus.. vertex is (-3, -2) a = 2 so the equation of the directrix is x = -2.... I'll let you figure the focus... just for something to do

OpenStudy (dumbcow):

solve for x x = (1/8)(y+2)^2 - 3 vertex: (-3, -2) directrix: x = -3 -1/(4a) a = 1/8 --> x = -3 -2 = -5 Focus: (-3 +2, -2) = (-1,-2)